Solve this pair of simultaneous equations by the elimination method:
4x + 2y = -8
x - 2y = -7

Answer by mananth(16946)

4.00 x + 2.00 y = -8.00 .............1
Total value
1.00 x -2.00 y = -7.00 .............2
Eliminate y
multiply (1)by 1.00
Multiply (2) by 1.00
4.00 x 2.00 y = -8.00
1.00 x -2.00 y = -7.00
Add the two equations
5.00 x = -15.00
/ 5.00
x = -3.00
plug value of x in (1)
4.00 x + 2.00 y = -8.00
-12.00 + 2.00 y = -8.00
2.00 y = -8.00 + 12.00
2.00 y = 4.00
y = 2.00
x= -3.00
y= 2.00